### v1

# Cospectral vertices, Walk-regular planar graphs and the echolocation problem

Abstract.
We study cospectral vertices on finite graphs in relation to the echolocation problem on Riemannian manifolds. First, We prove a computationally simple criterion to determine whether two vertices are cospectral. Then, we use this criterion in conjunction with a computer search to find minimal examples of various types of graphs on which cospectral but non-similar vertices exist, including minimal walk-regular non-vertex-transitive graphs, which turn out to be non-planar. Moreover, as our main result, we classify all finite walk-regular planar graphs, proving that such graphs must be vertex-transitive.
1. Introduction
Let $G$ be a finite graph with adjacency matrix $A$. Let $\phi_{1},\phi_{2},\ldots,\phi_{n}$ be an orthonormal basis of eigenfunctions (eigenvectors) of $A$ with respective eigenvalues $\lambda_{1},\lambda_{2},\ldots,\lambda_{n}$. We say two vertices $a$ and $b$ are cospectral if
(1.1) $\sum_{\lambda_{j}=\lambda}|\phi_{j}(a)|^{2}=\sum_{\lambda_{j}=\lambda}|\phi_{j}(b)|^{2}\qquad\text{ for all eigenvalues $\lambda$.}$
All graphs considered in this article will be finite, simple, undirected, and connected unless otherwise specified.
The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, ${}^{+}$19, EKL19, KSSW20]), especially due to their significant role in quantum information theory. It has been shown [God12] that for a perfect quantum state transfer to occur between two vertices in an interacting qubits network, it is necessary for these vertices to be cospectral.
We say two vertices $a$, $b$ are similar if there is an automorphism of $G$ mapping $a$ to $b$. Two similar vertices are clearly also cospectral. Recently, Kempton, Sinkovic, Smith, and Webb [KSSW20] show that two vertices are cospectral if and only if they are similar in a certain weaker sense. We are interested in the exceptional situation where two vertices are cospectral but not similar. We come to this problem from the smooth setting, where a graph is replaced with a Riemannian manifold. One can formulate a notion of cospectral points that reads exactly as (1.1), provided the adjacency matrix is replaced with the Laplace-Beltrami operator (see Remark 3.2). The question of whether or not there exists a pair of cospectral points that are not in the same orbit has a musical interpretation: Can one hear, up to symmetry, where a drum is struck? We refer to this question as the echolocation problem. The last two authors show that for most compact Riemannian manifolds, the answer is ‘yes’ [WX23]. So far, there are no known examples of connected manifolds where the answer is ‘no’.
In search of such negative examples, we turn to the graph-theoretic setting, where we might use a computer search. In fact, it is well-known that graphs with non-similar cospectral vertices exist. The first such example recorded in literature is a tree with 9 vertices observed by Schwenk [Sch73]. To find more such examples and study their properties, we must find a way to check when two vertices are cospectral that avoids floating point errors. The criterion (1.1) is not convenient in this way, so we need an alternative.
Our method builds off of work by Godsil and Smith [GS17], who show that two vertices $a$ and $b$ of a graph are cospectral with respect to the adjacent matrix $A$, if and only if $(A^{k})_{a,a}=(A^{k})_{b,b}$ for all $k\in\mathbb{N}.$ In other words, $a$ and $b$ are cospectral if, for all $k\in{\mathbb{N}}$, the number of walks of length $k$ starting and ending at $a$ is equal to the number of walks of length $k$ starting and ending at $b$. See also [CB23, Mon22] for results in other settings. When it comes to computation, this condition can be used to distinguish non-cospectral vertices but would require seemingly infinitely many operations to confirm that two vertices are cospectral. In this paper, we show that, in fact, one only needs to compute a finite number of powers of $A$ to check cospectrality for a finite graph.
Proposition 1.1.
Let $G$ be a graph (possibly with self-loops) with $n$ vertices and $A$ be the corresponding adjacent matrix. Then two vertices $a$ and $b$ are cospectral if and only if
(1.2) $(A^{k})_{a,a}=(A^{k})_{b,b},\quad\text{for all }0\leq k\leq n-1.$
Moreover, this condition is sharp in the sense that the upper limit $n-1$ cannot be made smaller.
We remark that, even though Godsil and Smith did not state it explicitly, the first part of Proposition 1.1 follows by the proof of [GS17, Lemma 2.1]. We shall include a proof of Proposition 1.1 in Section 2 for completeness. Then, we employ this condition in a computer search to find minimal examples of regular graphs with cospectral yet non-similar vertices in Section 3.
Theorem 1.2.
The smallest graph containing a pair of vertices that are cospectral but not similar has eight vertices.
Riemannian manifolds have some rigid structure that graphs, in general, lack. We impose a variety of kinds of regularity on our graphs to emulate this structure. We first search amongst the regular graphs for examples. This setting comes with the benefit that the adjacency matrix $A$ and the two standard symmetric normalizations of the Laplacian $L=D-A$ and $\mathcal{L}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ all have precisely the same eigenfunctions. Indeed, we find such examples, but only amongst graphs of ten or more vertices.
Theorem 1.3.
The smallest regular graph containing a pair of non-similar cospectral vertices has ten vertices.
Even more structured are walk-regular graphs [GM80], all of whose vertices are cospectral. A vertex $a$ has degree $(A^{2})_{a,a}$, and hence all walk-regular graphs are also regular. Seeking minimal examples amongst walk-regular graphs yields:
Theorem 1.4.
The smallest walk-regular graph containing a pair of non-similar vertices has twelve vertices.
Smooth analogs of walk-regular graphs were studied by Wang and the last two authors in [WWX23]. There we show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface. The assumption that our manifold is two-dimensional is essential to the proof, and it is yet unknown if there are any higher-dimensional manifolds that are walk-regular but not homogeneous. In addition, none of the walk-regular graphs we found are planar. This leads us to consider planar walk-regular graphs.
We are able to prove the following discrete analog of the result of [WWX23].
Theorem 1.5.
Every walk-regular planar graph with degree $d\geq 3$ is isomorphic to the polyhedral graphs of one of the following:
  • 5 regular polyhedrons, also known as Platonic solids: tetrahedron, hexahedron, octahedron, dodecahedron, and icosahedron;
  • 13 semi-regular polyhedrons, also known as Archimedean solids, pictured in Figure 1;
  • the $m$-gonal prisms for $m=3$ and $m\geq 5$; and
  • the $m$-gonal antiprisms for $m\geq 4$.
Consequently, every walk-regular planar graph is vertex-transitive.
A direct consequence of the classical theorem of Mani [Man71] says that all planar vertex-transitive finite graphs with degree at least 3 are exactly those listed above. Herbert and Wilfried [FI79] employed modified Schläfli symbol to give a direct proof of this fact. Theorem 1.5 is a refinement of this result of Mani and Herbert–Wilfried, as walk-regularity is, in general, strictly weaker than vertex-transitivity. The modified Schläfli symbol again plays a crucial role in our proof of Theorem 1.5.
Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12].
Organization of the paper
In Section 2, we prove Proposition 1.1. In Section 3, we give a short overview of our computer search and present a few of our minimal examples. The code in full is available on GitHub along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation). Section 4 is dedicated to the proof of Theorem 1.5.
Acknowledgements
Xi was supported by the National Key Research and Development Program of China No. 2022YFA1007200 and NSF China Grant No. 12171424. Kong was supported by NSF China Grant No. 12301109 and the Fundamental Research Funds for the Central Universities of China No. YJ202343. Wyman was supported by NSF grant DMS-2204397. The second listed author extends a thanks to Richard Lange for his very patient guidance through the process of setting up a GitHub repository.
2. Closed walks of bounded length determine if two vertices are cospectral
Let $A$ be the adjacency matrix of some graph $G$ on $n$ vertices, or any real-symmetric $n\times n$ matrix for that matter. Let $\phi_{1},\ldots,\phi_{n}$ be a real orthonormal eigenbasis for $A$ with respective eigenvalues $\lambda_{1},\ldots,\lambda_{n}$. For each vertex $a$, consider the measure
$\mu_{a}=\sum_{j}|\phi_{j}(a)|^{2}\delta_{\lambda_{j}}$
on ${\mathbb{R}}$. Observe that two vertices $a$ and $b$ are cospectral if and only if $\mu_{a}=\mu_{b}$.
Let $m\leq n$ denote the count of distinct eigenvalues and note each measure $\mu_{a}$ is determined by its pairing with degree $m-1$ polynomials. In particular, the measure $\mu_{a}$ is uniquely determined by its pairings
[…]
with the standard basis $\{1,t,\ldots,t^{m-1}\}$ for the degree $m-1$ polynomials. We can arrive at the same set of quantities from a different direction, namely
$A_{a,a}^{k}=\sum_{j,\ell}\phi_{j}(a)(\phi_{j}^{t}A^{k}\phi_{\ell})\phi_{\ell}(a)=\sum_{j}|\phi_{j}(a)|^{2}\lambda_{j}^{k}.$
In summary, we have shown:
Lemma 2.1.
Let $A$ be a real-symmetric $n\times n$ matrix. Then, $\mu_{a}=\mu_{b}$ if and only if
$A^{k}_{a,a}=A^{k}_{b,b}\qquad\text{ for all $k=0,1,\ldots,n-1$.}$
We note that the number $n-1$ of counts is sharp, since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
$A^{k}_{\ell,\ell}=A^{k}_{\ell+1,\ell+1}\quad\text{for all}~{}k=0,1,\ldots,n-2\quad\text{but}\quad A^{n-1}_{\ell,\ell}<A^{n-1}_{\ell+1,\ell+1}.$
3. Minimal examples
An exhaustive list of minimal examples of graphs with cospectral, non-similar vertices under regularity conditions of varying strength can be found on GitHub (https://github.com/ewy-man/graph-echolocation) along with the code used to generate them.
The code is written in the Julia language (see [BEKS17, ${}^{+}$21]) and contains a number of tools:
  1. (1)
    A few functions that generate exhaustive lists of non-isomorphic graphs on a specified number of vertices. An optional filter may be passed to these functions to narrow down and speed up the search.
  2. (2)
    A function which returns the orbits of the action of the isomorphism group on the vertices.
  3. (3)
    A function that determines the cospectrality of the vertices via Proposition 1.1.
  4. (4)
    A number of other convenience functions, such as those that determine if a graph is walk-regular, or if there exist non-similar cospectral pairs.
Once loaded into the Julia REPL, these functions can be called to generate the minimal examples below and to check that they are indeed minimal. The first listed operation is the most time-intensive, but typically does not take more than a minute to run on a personal computer.
3.1. Minimal graphs with non-similar cospectral vertices
The smallest graphs that contain non-similar cospectral pairs of vertices have eight vertices. Of all 12346 non-isomorphic graphs on 8 vertices, 126 of them contain non-similar cospectral vertices. Four examples are depicted in Figure 2. Furthermore, all of the 126 minimal examples are connected and none of them are trees.
Figure 2. Four examples of graphs with eight vertices with a pair of non-similar cospectral vertices. The non-similar cospectral pairs are marked in red.
3.2. Minimal regular graphs with non-similar cospectral vertices
The smallest regular graphs which contain non-similar cospectral pairs of vertices have ten vertices, and of those only those of degree 3, 4, 5, and 6 have minimal examples.
$d$ number of $d$-regular examples number of $d$-regular graphs
$3$ 3 21
$4$ 22 60
$5$ 22 60
$6$ 3 21
The symmetry in the table is not a coincidence.
Proposition 3.1.
Let $G=(V,E)$ be a finite graph and let $G^{\prime}=(V,E^{c})$ denote its edge complement.
  1. (1)
    $G$ and $G^{\prime}$ have the same graph isomorphism group.
  2. (2)
    If $G$ and $G^{\prime}$ are both connected, then two vertices are cospectral in $G$ if and only if they are cospectral in $G^{\prime}$.
The proposition explains the symmetry we see in the table. Indeed, all of the minimal examples counted by the table here happen to be connected. Each degree $3$ example’s edge complement is a minimal degree $6$ example, and similarly for degree $4$ and $5$ examples.
Remark 3.2.
Given a $d$-regular graph $G$ with adjacency matrix $A$, its graph Laplacian is written
$L=dI-A.$
As such, $A$ and $L$ share precisely the same eigenfunctions. Hence, two vertices $a$ and $b$ of a $d$-regular graph are cospectral if and only if
$\sum_{\lambda_{j}=\lambda}|\phi_{j}(a)|^{2}=\sum_{\lambda_{j}=\lambda}|\phi_{j}(b)|^{2}$
where $\phi_{1},\ldots,\phi_{n}$ is an orthonormal Laplace eigenbasis with $L\phi_{j}=\lambda_{j}\phi_{j}$.
Proof.
Part (1) is obvious and well-known. For part (2), note the Laplacian $L^{\prime}$ on $G^{\prime}$ can be written in terms of the Laplacian $L$ on $G$ as
$L^{\prime}=nI-\mathbf{1}-L,$
where $\mathbf{1}$ is the $n\times n$ matrix with a one in every entry. It suffices to show that an eigenfunction $\phi$ for $L$ is also an eigenfunction for $L^{\prime}$. Since $G$ and $G^{\prime}$ are both connected, their trivial eigenspaces are both spanned by the constant function, so it suffices to show $L$ and $L^{\prime}$ share nonconstant eigenfunctions. If $\phi$ is an eigenfunction of $L$ with eigenvalue $\lambda>0$, it is orthogonal to the constant eigenspace and hence lies in the kernel of $\mathbf{1}$. Hence,
$L^{\prime}\phi=(n-\lambda)\phi.$
The reverse argument is identical. ∎
We remark that, as a consequence, the edge complement of a walk-regular graph is again walk-regular. This is Theorem 4.2 in the work of Godsil and Mckay [GM80] on walk-regular graphs.
3.3. Minimal walk-regular, non-vertex-transitive graphs
There are no walk-regular, non-vertex-transitive graphs which have less than 12 vertices. There are exactly four minimal examples with twelve vertices, one each of degree 4, 5, 6, and 7. Again, these minimal examples are connected, and so can be put into edge-complement pairs; the examples of degrees $4$ and $7$ are edge complements, and so are the examples of $5$ and $6$. The degree $4$ and $5$ examples are depicted in Figure 3.
Refer to caption      Refer to caption
Figure 3. These graphs and their edge-complements are the only walk-regular, non-vertex-transitive graphs on $12$ vertices. Red vertices belong to one orbit and blue vertices belong to the other.
4. Walk-regular planar graphs
Here we will set up the tools we will need in order to classify the planar walk-regular graphs and prove Theorem 1.5.
[…]
$r(v)=(r_{1}(v),r_{2}(v),\ldots,r_{d}(v)),$
where $r_{i}(v)$ is the number of edges of $f_{i}(v)$. The first face $f_{1}(v)$, as well as the labeling direction, are chosen to minimize the label $r(v)$ lexicographically so that it is uniquely defined. In particular, $3\leq r_{1}(v)=\min_{1\leq i\leq d}r_{i}(v)$. With such a label, we count the total number $f$ of faces in the graph by
$f=\sum_{v}p(v),\qquad\text{where}\ \ p(v):=\sum_{i=1}^{d}\frac{1}{r_{i}(v)}.$
Using Euler’s formula $n+f-e=2$ and $e=nd/2$, we get
(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}$
We let $k(v)=\#\{1\leq i\leq d:r_{i}(v)=r_{1}(v)\}$. An easy observation is
Lemma 4.1.
In a walk-regular planar graph with $d\geq 3$,
[…]
Proof.
First we let $r_{1}=\min_{v}r_{1}(v)$. Then it follows from (4.1) that
$\frac{d}{r_{1}}\geq\frac{d-2}{2}+\frac{2}{n}.$
Thus $r_{1}\in\{3,4,5\}$ for $d=3$, and $r_{1}=3$ for $d=4$ or $5$. Let $v_{1}$ be a vertex such that $r_{1}(v_{1})=r_{1}$. Suppose that there is a vertex $v_{2}$ that satisfies $r_{1}(v_{2})>r_{1}$. If $r_{1}\in\{3,5\}$, we have
$A^{r_{1}}_{v_{2},v_{2}}=0<2k(v_{1})=A^{r_{1}}_{v_{1},v_{1}}.$
If $r_{1}=4$, then $d=3$ and
$A^{4}_{v_{2},v_{2}}=15<15+2k(v_{1})=A^{4}_{v_{1},v_{1}}.$
In any case, $v_{1}$ and $v_{2}$ are not cospectral. Hence $r_{1}(v)\equiv r_{1}$ provided that the graph is walk-regular. By a similar closed walk counting argument, we conclude that
$A_{v,v}^{r_{1}}-A_{v^{\prime},v^{\prime}}^{r_{1}}=2k(v)-2k(v^{\prime})$
for any pair of vertices $v,v^{\prime}$, and thus $k(v)\equiv k$ by walk-regularity. ∎
In order to prove Theorem 1.5, we shall prove the following lemma.
Lemma 4.2.
For any walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
(4.2) $r(v)\equiv(r_{1},r_{2},\cdots,r_{d})\qquad\textit{for all vertices}~{}v.$
However, the proof of this lemma appears to be quite delicate. The 3-regular planar graph, as illustrated in Figure 5, features cospectral vertices $v_{1}$ and $v_{2}$ with $r(v_{1})=(3,5,5)$ but $r(v_{2})=(3,5,6)$. Despite vertex $v_{1}$ having two neighboring pentagons and $v_{2}$ only one, the adjacent triangle $\{v_{5},v_{7},v_{9}\}$ contributes two additional closed walks of length 5 to $v_{2}$. Moreover, this graph is ‘almost’ walk-regular, with all vertices except $v_{4}$ being cospectral. Therefore, establishing (4.2) succinctly appears challenging. Instead, we justify (4.2) in each individual case.
We remark that, once we prove Lemma 4.2, it is routine to check that Theorem 1.5 holds in each case. Indeed, the numbers of vertices, edges, and faces are determined by Lemma 4.2 and Euler’s identity. The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer polygon. We illustrate this fact for the most complicated case when $r(v)\equiv(3,4,5,4)$ in Figure 4. The only exception is the case when $r(v)\equiv(3,4,4,4)$, for which one runs into two possible reconstructions, the polyhedral graphs of the $(3,4,4,4)$-solid and twisted $(3,4,4,4)$-solid. However, the twisted one is not walk-regular. See Case (13) in the proof.
Figure 4. Step by step reconstruction of the $(3,4,5,4)$-solid.
Figure 5. A 3-regular planar graph on 10 vertices.
Let $v_{1}$ be a vertex such that $p(v_{1})=\max_{v}p(v)$. In view of Lemma 4.1, to prove Lemma 4.2, we only need to discuss the sixteen cases enumerated below. For the convenience of the reader, we have also included an exhaustive list of the possible walk-regular planar graphs belonging to each case. From now on and throughout the paper, for the purpose of simplicity, we shall not distinguish between a solid and its polyhedral graph.
Case $d,r_{1},k$ Solids
(1) 3, 3, 3 Tetrahedron
(2) 3, 3, 2
(3) 3, 3, 1 Triangular prism, $(3,6,6)$-, $(3,8,8)$-, and $(3,10,10)$-solids
(4) 3, 4, 3 Hexahedron
(5) 3, 4, 2 $m$-gonal prism, $m\geq 5$
(6) 3, 4, 1 $(4,6,6)$-, $(4,6,8)$-, and $(4,6,10)$-solids
(7) 3, 5, 3 Dodecahedron
(8) 3, 5, 2
(9) 3, 5, 1 $(5,6,6)$-solid
(10) 4, 3, 4 Octahedron
(11) 4, 3, 3 $m$-gonal antiprism, $m\geq 4$
(12) 4, 3, 2 $(3,4,3,4)$- and $(3,5,3,5)$-solids
(13) 4, 3, 1 $(3,4,4,4)$- and $(3,4,5,4)$-solids
(14) 5, 3, 5 Icosahedron
(15) 5, 3, 4 $(3,3,3,3,4)$- and $(3,3,3,3,5)$-solids.
(16) 5, 3, $\leq$ 3
Of these cases, (2) and (8) are simply impossible to realize. It is also impossible to realize case (16), since
$p(v)\leq\frac{1}{3}+\frac{1}{3}+\frac{1}{3}+\frac{1}{4}+\frac{1}{4}=\frac{3}{2}$
violates (4.1). One quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), (7), (10), and (14) must be constant.
Next, we focus on cases (5), (11), (9), and (15), which are easier to resolve. The remaining cases (3), (6), (12), and (13) are more involved. We will treat these cases after.
Case (5)
($d=3,r_{1}=4,k=2$) For each $m=r_{3}(v_{1})\geq 5$, the face $f_{3}(v_{1})$ surrounded by $m$ quadrangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
Case (11)
($d=4,r_{1}=3,k=3$) For each $m=r_{4}(v_{1})\geq 4$, the face $f_{4}(v_{1})$ surrounded by $m$ triangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
Starting from the next case, we often need to count the closed walks of length $m$ for some $m\geq 4$ and two adjacent vertices $v,v^{\prime}$ with $r(v)=r(v_{1})\neq r(v^{\prime})$. Our strategy of counting is to consider only the $\lfloor\frac{m}{2}\rfloor$-neighborhood of $\{v,v^{\prime}\}$, which turns out to be determined by the values of $d,r_{1},k$ and $r(v_{1})$, with at most two possibilities.
Case (9)
($d=3,r_{1}=5,k=1$) Each face with $m\geq 6$ edges must be surrounded by exact $m/2$ pentagonal faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), the only possibility is $r(v_{1})=(5,6,6)$. Suppose $r(v)\not\equiv(5,6,6)$. Then there exists a pentagonal face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$. The $3$-neighborhood of $\{v,v^{\prime}\}$ is unique, with 23 vertices (Figure 6). This gives
$A_{v,v}^{6}=91>89=A^{6}_{v^{\prime},v^{\prime}}$
Hence $r(v)\equiv(5,6,6)$, the graph is isomorphic to the $(5,6,6)$-solid.
Figure 6. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$.
Case (15)
($d=5,r_{1}=3,k=4$) As $p(v_{1})>\frac{3}{2}$ by (4.1), $r_{5}(v_{1})=4$ or $5$. When $r_{5}(v_{1})=4$, we suppose $r_{5}(v)\not\equiv 4$. Then there exist two adjacent vertices $v,v^{\prime}$ such that $r_{5}(v)=4<r_{5}(v^{\prime})$. Counting closed walks of length $4$ in the $2$-neighborhood of $\{v,v^{\prime}\}$ (Figure 7), we have
$A_{v,v}^{4}=59>57=A_{v^{\prime},v^{\prime}}^{4}.$
Hence $r(v)\equiv(3,3,3,3,4)$, the graph is isomorphic to the $(3,3,3,3,4)$-solid.
Figure 7. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,4)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>4$.
In the case that $r_{5}(v_{1})=5$, $r_{5}(v)\not\equiv 5$, there must exist two adjacent vertices $v,v^{\prime}$ with $r_{5}(v)=5<r_{5}(v^{\prime})$. Counting closed walks of length 5 in the 2-neighborhood of $\{v,v^{\prime}\}$ (Figure 8) gives
$A_{v,v}^{5}=182>180=A_{v^{\prime},v^{\prime}}^{5}.$
Hence $r(v)\equiv(3,3,3,3,5)$, the graph is isomorphic to the $(3,3,3,3,5)$-solid.
Figure 8. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,5)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>5$.
We now move on to the more involved cases.
Case (3)
($d=3,r_{1}=3,k=1$) Denote the vertex set of the triangular face $f_{1}(v_{1})$ by $\{v_{1},v_{2},v_{3}\}$, where $v_{2}$ does not belong to the face $f_{2}(v_{1})$. Note that every face with $m\geq 4$ edges should be surrounded by exact $m/2$ triangular faces, and thus $m$ must be even. Suppose $r_{2}(v_{1})>10$. Then
$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
which violates (4.1). Hence $r_{2}(v_{1})=4,6,8$ or $10$.
Sub-case: $r_{2}(v_{1})=4$
If $r_{3}(v_{1})>4$, it follows that $r_{2}(v_{2})>4$ and
$A_{v_{1},v_{1}}^{4}=17>15=A_{v_{2},v_{2}}^{4}.$
Hence the walk-regularity implies $r_{3}(v_{1})=4$. In this case, the only possibility is $r(v)\equiv(3,4,4)$, and the graph is isomorphic to the triangular prism.
Sub-case: $r_{2}(v_{1})=6$
Suppose $r_{3}(v_{1})>6$. Then $r_{2}(v_{2})>6$, otherwise we have $r(v_{3})=(3,6,6)$ that violates the minimality of $p(v_{1})$. Counting the closed walks with length $6$ in the $3$-neighborhood of $\{v_{1},v_{2}\}$ (Figure 9)
$A_{v_{1},v_{1}}^{6}=91>89=A_{v_{2},v_{2}}^{6}.$
Thus $r(v_{1})=(3,6,6)$ by the walk-regularity.
Figure 9. 3-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,6,m)$ with $m>6$.
Now we claim that $r(v)\equiv(3,6,6)$. Indeed, whenever this is false, there exists a triangular face with vertex set $\{v,v^{\prime},v^{\prime\prime}\}$ such that $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$ (Figure 10), which gives
$A_{v,v}^{6}=93>91=A_{v^{\prime},v^{\prime}}^{6}.$
Hence $r(v)\equiv(3,6,6)$, and the graph is isomorphic to the $(3,6,6)$-solid.
Figure 10. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$.
Sub-case: $r_{2}(v_{1})=8$
As in the previous sub-case, $r_{3}(v_{1})>8$ implies $r_{2}(v_{2})>8$, and a counting in the $4$-neighborhood of the pair $\{v_{1},v_{2}\}$ (Figure 11) provides
$A_{v_{1},v_{1}}^{8}=593>591=A_{v_{2},v_{2}}^{8},$
thus $r(v_{1})=(3,8,8)$.
Figure 11. 4-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,8,m)$ with $m>8$.
If $r(v)\equiv(3,8,8)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$ (Figure 12), and
$A_{v,v}^{8}=595>593=A_{v^{\prime},v^{\prime}}^{8}.$
This shows that $r(v)\equiv(3,8,8)$, and the graph is isomorphic to the $(3,8,8)$-solid.
Figure 12. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$.
Sub-case: $r_{2}(v_{1})=10$
Again, $r_{3}(v_{1})>10$ implies $r_{2}(v_{2})>10$, and a counting in the 5-neighborhood of $\{v_{1},v_{2}\}$ (Figure 13) gives
$A_{v_{1},v_{1}}^{10}=4225>4223=A_{v_{2},v_{2}}^{10},$
thus $r(v_{1})=(3,10,10)$.
Figure 13. 5-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,10,m)$ with $m>10$.
If $r(v)\equiv(3,10,10)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$ (Figure 14), and
$A_{v,v}^{10}=4227>4225=A_{v^{\prime},v^{\prime}}^{10}.$
This shows that $r(v)\equiv(3,10,10)$, and the graph is isomorphic to the $(3,10,10)$-solid.
Figure 14. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$.
Case (6)
($d=3,r_{1}=4,k=1$) Each face with $m\geq 5$ should be surrounded by exact $m/2$ quadrangular faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), we must have $r_{2}(v_{1})=6$ and $r_{3}(v_{1})=6,8$ or $10$.
Sub-case: $r_{3}(v_{1})=6$
Suppose $r(v)\not\equiv(4,6,6)$. Then there exists a quadrangular face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$ (Figure 15). By counting closed walks we obtain
$A_{v,v}^{6}=115>113=A_{v^{\prime},v^{\prime}}^{6}.$
Thus $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
Figure 15. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$.
Sub-case: $r_{3}(v_{1})=8$
For this we claim that for every vertex $v$ with $r(v)=(4,6,8)$, its adjacent vertex $v^{\prime}$ outside the hexagonal face $f_{2}(v)$ also has $r(v^{\prime})=(4,6,8)$. Indeed, if this is violated by some adjacent pair $v,v^{\prime}$ (Figure 16), then
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This verifies the claim.
Figure 16. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r_{2}(v^{\prime})>6$.
Now suppose $r(v)\not\equiv(4,6,8)$. Then by the claim there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$ (Figure 17). A counting of closed walks with length $8$ gives
$A_{v,v}^{8}=811>809=A_{v^{\prime},v^{\prime}}^{8}.$
Hence $r(v)\equiv(4,6,8)$, and the graph is isomorphic to the $(4,6,8)$-solid.
Figure 17. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$.
Sub-case: $r_{3}(v_{1})=10$
Note that any two adjacent vertices $v,v^{\prime}$ with $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$ (Figure 18) satisfy
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This shows that any vertex that is adjacent to some vertex with label $(4,6,10)$ should lie on a hexagonal face.
Figure 18. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$.
Moreover, if $r(v)\not\equiv(4,6,10)$, then there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$ (Figure 19), which give
$A_{v,v}^{10}=6065>6063=A_{v^{\prime},v^{\prime}}^{10}.$
Hence $r(v)\equiv(4,6,10)$, and the graph is isomorphic to the $(4,6,10)$-solid.
Figure 19. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$.
Case (12)
($d=4,r_{1}=3,k=2$) We start with a lemma.
Lemma 4.3.
Let $G$ be $d$-regular with $d\geq 3$. Then,
$A_{v_{i},v_{i}}^{4}-(2d^{2}-d)=2(\#\text{ of $4$-cycles through $v_{i}$}).$
Proof.
Every closed walk of length $4$ is either contractible or traces a $4$-cycle in one of two directions, and there are exactly $d(d-1)+d^{2}$ contractible closed walks of length $4$. ∎
We understand each vertex is adjacent to two triangles and two other faces, say of degree $4\leq m\leq m^{\prime}$ respectively, both larger than $3$. We subdivide the argument into four sub-cases where there are zero, one, two, or more than two $4$-cycles through each vertex.
Sub-case: Zero 4-cycles
There are no squares, and no two triangles share an edge. Such graphs have plenty of structure. Firstly, each face of $\geq 5$ sides is adjacent only to triangular faces. Furthermore, each such face is, through each vertex, opposed to another face of $\geq 5$ sides. We first establish that, in this situation,
$A_{v_{i},v_{i}}^{5}-60=2\times(\#\text{ of pentagons incident to $v_{i}$})$
Our job is now to eliminate the cases where
$r(v)=(3,5,3,6)\text{ or }(3,5,3,7)$
for each $v$, and we will be left with $r(v)\equiv(3,5,3,5)$. In fact, if we have either case above for one vertex, we have it for all of them. We can construct a $3$-neighborhood of the vertex $v$ (Figure 20) with either a hexagon or a face of $\geq 7$ sides and obtain
$A_{v,v}^{6}=270+\begin{cases}2&r(v)=(3,5,3,6)\\ 0&r(v)=(3,5,3,7).\end{cases}$
Figure 20. Two possible 3-neighborhood of $v$.
If we attempt to construct either graph with either $r\equiv(3,5,3,6)$ or $r\equiv(3,5,3,7)$, we quickly run into absurdities. Hence, we are left with the $(3,5,3,5)$ solid.
Sub-case: one $4$-cycle
First, we rule out the possibility that such a graph has a square face. If we start with a square face, we obtain a graph resembling that in Figure 21, which we will call a “bee” since it looks like it has a head with two antennae, a thorax with two wings, and an abdomen.
Figure 21. Our “bee” graph.
Note, one or both of the antennae may attach to a wing to form a pentagon. Running some software, we find that, for any vertex $v$ of the square,
$A_{v,v}^{5}=\begin{cases}68&\text{ if no pentagon is incident to $v$}\\ 70&\text{ if the remaining face incident to $v$ is a pentagon}.\end{cases}$
On the other hand, the vertex $v^{\prime}$ at the “neck” of the bee satisfies
$A_{v^{\prime},v^{\prime}}^{5}=62+2\times(\#\text{ of pentagons incident to $v^{\prime}$})\leq 66,$
and so we have $A_{v^{\prime},v^{\prime}}^{5}<A_{v,v}^{5}$, a contradiction.
We have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces. Since there are no squares, there must be at least one pentagon. The graph self-destructs upon any attempt to draw a $1$-neighborhood of the pentagon.
Sub-case: two $4$-cycles
There are a few situations we must rule out. We first establish that each vertex must be incident to a square face. Suppose $v$ is not adjacent to a square face. Then, $r(v)=(3,3,m,m^{\prime})$ or $(3,m,3,m^{\prime})$, where $m,m^{\prime}\geq 5$. The first is incident to only one $4$-cycle, so is ruled out. We try to draw the graph starting at this vertex and obtain the one in Figure 22, which forces a vertex to have at least three $4$-cycles.
Figure 22. The left-center vertex is forced to be incident to three $4$-cycles.
With this in hand, any construction of a graph with a vertex with the label $(3,3,4,m)$ for $m\geq 5$ fails after drawing a few faces. Hence, every vertex of our graph must have label $(3,4,3,4)$, and we obtain the $(3,4,3,4)$-solid.
Sub-case: More than two $4$-cycles
The only way a vertex can be incident to four $4$-cycles is if its $2$-neighborhood is that in Figure 23. However, this vertex’s neighbors are only ever incident to three $4$-cycles.
Figure 23. $2$-neighborhood of a vertex incident to four $4$-cycles.
Hence, all that remains is to exclude the case where the vertices are incident to exactly three $4$-cycles.
A similar argument as the one above excludes the case where $v$ is incident to a face of $5$ or more sides. Hence, for each vertex $v$, we have either $r(v)=(3,4,3,4)$ or $r(v)=(3,3,4,4)$. There is a unique graph with these properties, depicted in Figure 24.
Figure 24. The graph with exactly three $4$-cycles through each vertex.
This graph, however, is not walk-regular. The three innermost and outermost vertices all have $74$ returning walks of length $5$, while the other vertices have $76$.
Case (13)
($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to the number of 4-cycles through each vertex. Since there is exactly 1 triangle face adjacent to each vertex, the number of 4-cycles is determined by the number of adjacent squares. Therefore, each vertex must be adjacent to the same number of squares. By (4.1), we must have $p(v_{1})>1$, and thus there must be at least two squares adjacent to each vertex.
Sub-case: three squares at each vertex
This case is straightforward since we already know that $r(v)\equiv(3,4,4,4)$. We remark that, in this case, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph. However, there are vertices on the twisted $(3,4,4,4)$ solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out. See Figure 25.
Figure 25. (3,4,4,4)-solid and twisted (3,4,4,4)-solid.
Sub-case: two squares at each vertex
In this case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$. We first consider the case $r(v_{1})=(3,4,4,5)$. One can try to draw such a graph starting from $v_{1}$ and end up with a neighborhood of the shape as in Figure 26.
Figure 26. $r(v_{1})=(3,4,4,5)$
Here, vertex $v_{1}$ is what we started with. We claim that the unknown polygon adjacent to vertex $v_{3}$ must be a pentagon. Then a similar argument will show that all outer polygons must be pentagons. To do so, we count the closed walks of length 5 for $v_{1}$ and for $v_{3}$. Non-trivial ones are either a pentagon or a triangle attached to a square, which we shall call a “house”. $v_{1}$ and $v_{3}$ always share the same number of “houses”, and we are done with the claim.
Next, we claim that we will get a contradiction if the center polygon is not a pentagon. In this case, for $v_{1}$ and $v_{2}$ to have the same number of closed walks of length 5, we cannot create more “houses” in the outer layer for vertex $v_{1}$. This forces us to get 3 squares in a row. After drawing all the squares we arrive at Figure 27, which is impossible since there are vertices adjacent to more than 2 triangles.
Figure 27. An impossible graph when the center is not a pentagon.
Therefore, we conclude that we must have a pentagon at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 28.
Figure 28. 3-neighborhood of $v_{1}$ with $r(v_{1})=(3,4,4,5).$
We remark that this is exactly part of the twisted $(3,4,5,4)$-solid. To get a contradiction, we count the number of closed walks of length 6 passing through $v_{1}$ and through $v_{2}$. The non-trivial shapes that contribute to the number of closed walks of length 6 are graphed in Figure 29.
Figure 29. Possible non-trivial shapes contributing to the number of closed walks of length 6.
One can see that $v_{1}$ and $v_{2}$ share the same count for all shapes except for the double-square, for which $v_{1}$ has 1 but $v_{2}$ has 2, a contradiction. Therefore, there is no vertex $v$ satisfying $r(v)=(3,4,4,5)$ in this sub-case, and $v_{1}$ satisfies $r(v_{1})=(3,4,5,4)$. Again by counting closed walks of length 5, one can easily show that all vertices adjacent to $v_{1}$ must be adjacent to a pentagon as well, and as a consequence, $r(v)\equiv(3,4,5,4)$. The proof is complete.